A roulette wheel consists of 18 black slots, 18 red slots, and 2 green slots. If a gambler bets $10 on red, what is the gamblers expected gain or loss?
1
Expert's answer
2021-09-15T02:57:58-0400

Let the winning amount be w.

There are 18 red slots out of 18 + 18 + 2 = 38.

Probability that the bet on red is a success "= \frac{18}{38}"

Probability that the bet on red is not a success "= 1 - \frac{18}{38} = \frac{20}{38}"

=> Expected value "= w \times \frac{18}{38} - \$10 \times \frac{20}{38}"

If w = $10 for instance, expected loss "= \$10 \times \frac{18}{38} - \$10 \times \frac{20}{38}"

= -0.5263$.